# Mitigating the Accuracy-Robustness Trade-off via Multi-Teacher Adversarial Distillation

### v1

@@ Abstract
Adversarial training is a practical approach for improving the robustness of deep neural networks against adversarial attacks. Although bringing reliable robustness, the performance toward clean examples is negatively affected after adversarial training, which means a trade-off exists between accuracy and robustness. Recently, some studies have tried to use knowledge distillation methods in adversarial training, achieving competitive performance in improving the robustness but the accuracy for clean samples is still limited. In this paper, to mitigate the accuracy-robustness trade-off, we introduce the Multi-Teacher Adversarial Robustness Distillation (MTARD) to guide the model’s adversarial training process by applying a strong clean teacher and a strong robust teacher to handle the clean examples and adversarial examples, respectively. During the optimization process, to ensure that different teachers show similar knowledge scales, we design the Entropy-Based Balance algorithm to adjust the teacher’s temperature and keep the teachers’ information entropy consistent. Besides, to ensure that the student has a relatively consistent learning speed from multiple teachers, we propose the Normalization Loss Balance algorithm to adjust the learning weights of different types of knowledge. A series of experiments conducted on public datasets demonstrate that MTARD outperforms the state-of-the-art adversarial training and distillation methods against various adversarial attacks.
@@ 1 Introduction
Deep Neural Networks (DNNs) have become powerful tools for solving complex real-world learning problems, such as image classification [13], face recognition [34], and natural language processing [29]. However, Szegedy et al. [32] demonstrate that DNNs are vulnerable to adversarial attacks with imperceptible adversarial perturbations on input, which causes wrong predictions of DNNs. To defend against adversarial attacks, adversarial training is proposed and has shown its effectiveness in obtaining adversarial robust DNNs [35, 6, 25]. While improving the robustness of DNNs, a negative impact exists on the model’s accuracy on clean samples. Some methods [45, 26] are proposed to alleviate the trade-off between accuracy and robustness. However, this phenomenon still exists and needs to be further explored.
@@
Recently, to further enhance the robustness of small DNNs, Knowledge Distillation [15] is applied as a powerful tool in adversarial training. Knowledge Distillation can transfer knowledge from strong robust models to the student model, which utilizes the predictions from the teacher models as label information to guide the student model [9, 48, 49]. These methods are called adversarial robustness distillation. Compared with one-hot ground truth labels, the teacher’s predicted labels can not only retain the correctness of the target label but also reflect richer knowledge information of the non-target label. Several studies attempt to explain why knowledge distillation is effective through various views, e.g., Re-weighting [8], Privileged Information [24], and Label-smoothing [43]. Recently, Li et al. [20] argue that correctness, smooth regularization, and class discriminability in teacher’s predicted labels can make up the knowledge of the teacher’s predicted distribution, and that appropriate category discriminability plays an important role in improving the effectiveness of knowledge distillation. While achieving impressive robust performance, we are curious if adversarial robustness distillation could address the trade-off between accuracy and robustness.
@@
Generally speaking, if we want to improve both accuracy and robustness through knowledge distillation, the provided teacher model is supposed to perform well in both aspects. That is to say, the teacher model should equally provide correct guidance and rich knowledge information for clean and adversarial examples. However, due to the existing trade-off, a single teacher model is difficult to achieve the above goals. Therefore, a reasonable approach is to decouple adversarial knowledge distillation, i.e., using two teachers that are good at accuracy and robustness respectively (they can be called the clean teacher and the robust teacher) to guide the student in different knowledge types. Under this divide-and-conquer strategy, the student model guided by multiple teacher models can enhance its performance in both aspects theoretically.
@@
However, the actual implementation of this idea is challenging. Owing to the intrinsic trade-off between accuracy and robustness, the clean teacher and robust teacher push the student model towards opposite directions. This is obviously different from the current multi-teacher knowledge distillation methods [31, 42, 23], where their optimization goal of different teachers is consistent, i.e., all teachers push the student towards the good clean accuracy. The trade-off in our task leads to a difficult optimization mainly in two aspects: (1) The two teacher models should have similar knowledge scales themselves. When one teacher has too much knowledge scale, the student will be forced to obtain more knowledge from this teacher and relatively less knowledge from another, which directly results in an imbalance between accuracy and robustness. However, finding a clean teacher and a robust teacher with similar knowledge scales is not easy. Due to the different network architecture and training methods, a natural gap exists in knowledge scales between different teachers. Meanwhile, how to accurately measure teachers’ knowledge scales is also an unsolved problem. (2) The student’s learning ability from different types of knowledge should be balanced. More specifically, because two hugely different optimization directions exist in adversarial robustness distillation, the student may over-fit an ability that is easier to learn but ignore another due to different learning difficulties. Only the cooperation of the student and teachers can help the student obtain both strong accuracy and robustness.
@@
Based on the above discussions, in this paper, we propose the Multi-Teacher Adversarial Robustness Distillation (MTARD) to enhance both accuracy and robustness at the same time. During adversarial training, a clean teacher is to instruct the student for handling clean examples and a strong robust teacher is to instruct the student for handling adversarial examples. To meet the two challenges above, the novel Entropy-Based Balance and Normalization Loss Balance strategies are presented.
@@
For the second challenge, to maintain the relative equality of the student’s learning speed from multiple teachers, we propose a Normalization Loss Balance algorithm to control loss weight and balance the influence between the clean teacher and the robust teacher. Specifically, inspired by [4], we design a relative loss to measure the knowledge proportion that the student learns from different teachers. The different loss weights for the update of student parameters are dynamically adjusted to keep the relative loss consistent during the training process. Compared with the initial state, the student model will be forced to learn relatively equal knowledge scales from all the teacher models.
@@
  • We propose a novel framework: Multi-Teacher Adversarial Robustness Distillation (MTARD). MTARD decouples the adversarial knowledge distillation to obtain accuracy and robustness respectively. We apply a clean teacher and a robust teacher to bring both clean and robust knowledge to the student.
@@
  • We empirically verify the effectiveness of MTARD in improving performance. The Weighted Robust Accuracy (a measure to evaluate both clean and robust accuracy) of our MTARD-trained models improves significantly against a variety of attacks compared to state-of-the-art adversarial training and knowledge distillation methods. Especially for query-based black-box attacks, MTARD can enhance the Weighted Robust Accuracy by 5.56%, 4.78%, and 3.31% for MobileNet-v2 on CIFAR-10, CIFAR-100, and Tiny-ImageNet, respectively.
This journal paper is an extended version of our ECCV paper [47]. Compared with the conference version, we have made significant improvements and extensions in this version in the following aspects: (1) At the idea level, we discuss the trade-off problem in a more comprehensive view from both teachers and the student, and then point out that the difference in the two teachers’ knowledge scales will affect the trade-off between accuracy and robustness (Section 1). (2) At the method level, a multiple teacher’s knowledge adjustment mechanism named the Entropy-Based Balance algorithm is designed to implement the new idea (Section 3.2). (3) At the experimental level, we additionally train models on the Tiny-ImageNet dataset and compare our proposed method with more SOTA methods against more advanced attacks. In addition, we give more ablation studies to comprehensively test our method (Section 4). (4) At the writing level, the introduction, related works, methods, and experiments have been rewritten and refined.
The rest of the paper is organized as follows: We introduce the related work in Section 2. Section 3 proposes our MTARD to enhance both accuracy and robustness. The experiments are conducted in Section 4, and the conclusion is given in Section 5.
@@ 2.1 Adversarial Attacks
Since Szegedy et al. [32] propose that adversarial examples can mislead the deep neural network, lots of effective adversarial attack methods, such as the Fast Gradient Sign Method (FGSM) [10], Projected Gradient Descent Attack (PGD) [25], and Carlini and Wagner Attack (CW) [3] are proposed. Attack methods can be divided into white-box attacks and black-box attacks. White-box attacks know all the parameter information of the target model when generating adversarial examples, and black-box attacks know little or no information. In general, black-box attacks simulate the model gradient by repeatedly querying the target model (query-based attack) [1, 39, 36, 37, 38] or searching for an alternative model similar to the target model (transfer-based attack) [7, 16]. Recently, a strong ensemble adversarial attack method named AutoAttack (AA) [6] is proposed, which consists of four attack methods, including Auto-PGD (APGD), Difference of Logits Ratio (DLR) attack, FAB-Attack [5], and the black-box Square Attack [1].
@@ 2.2 Adversarial Training
Much work has been proposed to further improve the robustness of adversarial training. Zhang et al. [45] attempt to reduce the gap between accuracy and robustness by minimizing a Kullback–Leibler (KL) divergence loss (TRADES). Wang et al. [35] further improve performance through Misclassification-Aware Adversarial Training (MART). Pang et al. [26] propose to employ local equivariance to describe the ideal behavior of a robust model, which facilitates the reconciliation between accuracy and robustness (SCORE).
Previous work [45, 26] uses two indexes: clean and robust accuracy, as the metric to evaluate the model’s performance. Gurel et al. [12] propose a metric named Weighted Robust Accuracy to comprehensively measure the model’s performance, which is used in our experiments.
@@ 2.3 Knowledge Distillation
where $KL$ is Kullback–Leibler divergence loss, $CE$ represents the cross-entropy loss. $\tau$ is a temperature constant combined with softmax operation, and the prediction $p_{k}$ with temperature $\tau$ can be formulated as follows:
$p_{k}=\frac{exp(z_{k}(x)/\tau)}{\sum_{j=1}^{C}exp(z_{j}(x)/\tau)},$ (3)
where $z_{k}(x)$ denotes the $k$-th dimension output logits of model before softmax layer, and $p_{k}$ denotes the predicted possibility of $k$-th class.
@@
Some previous studies exist on multi-teacher knowledge distillation from different views: including response-based knowledge [18, 42], feature-based knowledge [23, 46], and relation-based knowledge [41, 40]. Previous methods often use multiple teachers for knowledge distillation for the same examples, and those teachers tend to optimize the student in similar directions. Different from previous methods, our multi-teacher adversarial robustness distillation targets two different kinds of examples. Since adversarial examples are generated to mislead the model, there is a huge gap between the optimization direction of clean examples and adversarial examples, which exist great difficulties compared with previous research.
@@ 2.4 Adversarial Robustness Distillation
To defend against the adversarial attack, the defense distillation [27] is proposed to use the knowledge distillation to defend against adversarial attack. However, this method is considered as “Obfuscated Gradients” [2] and can be broken by CW attack [3]. Then a series of methods [49, 9, 48] apply knowledge distillation to further enhance the robustness based on adversarial training [25]. ARD [9] proposes that using a strong robust model as the teacher model in adversarial training can achieve pretty robustness. IAD [48] performs adversarial knowledge distillation by compositing with unreliable teacher guidance and student introspection. RSLAD [49] uses the soft labels generated by the robust teacher model to produce adversarial examples, and further utilizes the robust labels to guide the training process of both clean examples and adversarial examples, which can effectively improve the robustness of the student.
The above methods focus on improving the robustness but ignore the performance of identifying clean examples. In our method, we consider both accuracy and robustness through clean and robust teachers, which comprehensively improves the performance of the student model.
@@ 3 Methodology
In this section, we propose MTARD to enhance both accuracy and robustness by applying the clean teacher and the robust teacher to guide the student; From the teachers’ teaching view, the Entropy-Based Balance algorithm is proposed to balance the teacher’s knowledge scale. From the student’s learning view, the Normalization Loss Balance algorithm is designed to balance the student’s learning speeds from different teachers.
Figure 1: The framework of our Multi-Teacher Adversarial Robustness Distillation (MTARD). In the training process, we first generate adversarial examples from the student. Then, with the guidance of the clean teacher and the robust teacher, the student is trained on clean examples and adversarial examples, respectively. Here, we apply the Entropy-Based Balance algorithm to adjust the teacher’s knowledge scales; In addition, we use the Normalization Loss Balance algorithm to balance the student’s knowledge learning speed from different teachers.
3.1 Multi-Teacher Adversarial Robustness Distillation
The training process in MTARD can be considered as a min-max optimization based on adversarial training. The whole framework can be viewed in Fig. 1. In the inner maximization, the adversarial examples are generated by the student based on the clean examples. In the outer minimization, the clean teacher and the robust teacher generate predicted labels to guide the student training process toward the clean examples and the adversarial examples, respectively. The optimization of the basic MTARD is defined as follows:

### v3

# Mitigating Accuracy-Robustness Trade-off via Balanced Multi-Teacher Adversarial Distillation

@@ Abstract
Adversarial Training is a practical approach for improving the robustness of deep neural networks against adversarial attacks. Although bringing reliable robustness, the performance towards clean examples is negatively affected after Adversarial Training, which means a trade-off exists between accuracy and robustness. Recently, some studies have tried to use knowledge distillation methods in Adversarial Training, achieving competitive performance in improving the robustness but the accuracy for clean samples is still limited. In this paper, to mitigate the accuracy-robustness trade-off, we introduce the Balanced Multi-Teacher Adversarial Robustness Distillation (B-MTARD) to guide the model’s Adversarial Training process by applying a strong clean teacher and a strong robust teacher to handle the clean examples and adversarial examples, respectively. During the optimization process, to ensure that different teachers show similar knowledge scales, we design the Entropy-Based Balance algorithm to adjust the teacher’s temperature and keep the teachers’ information entropy consistent. Besides, to ensure that the student has a relatively consistent learning speed from multiple teachers, we propose the Normalization Loss Balance algorithm to adjust the learning weights of different types of knowledge. A series of experiments conducted on three public datasets demonstrate that B-MTARD outperforms the state-of-the-art methods against various adversarial attacks.
@@ 1 Introduction
Deep Neural Networks (DNNs) have become powerful tools for solving complex real-world learning problems, such as image classification [17], face recognition [41], and natural language processing [35]. However, Szegedy et al. [39] demonstrate that DNNs are vulnerable to adversarial attacks with imperceptible adversarial perturbations on input, which causes wrong predictions of DNNs. To defend against adversarial attacks, Adversarial Training is proposed and has shown its effectiveness in obtaining adversarial robust DNNs [42, 10, 31]. While improving the robustness of DNNs, a negative impact exists on the model’s accuracy on clean samples. Some methods are proposed to alleviate the trade-off between accuracy and robustness from different perspectives, e.g., optimization [55, 42, 32, 50, 56] and extra data [1, 5, 19]. However, this phenomenon still exists and needs to be further explored.
@@
Recently, to further enhance the robustness of small DNNs, Knowledge Distillation [20] is applied as a powerful tool in Adversarial Training, which can transfer knowledge from strong robust models to the student model. It utilizes the teacher’s predictions as label information to guide the student model within the framework of Adversarial Training, including the optimization to generate the adversarial examples (Maximization process) and apply adversarial examples to train the student with the assistance of the teacher (Minimization process). These methods are called adversarial robustness distillation (ARD) [13, 60, 61]. Compared with one-hot ground truth labels, the teacher’s predicted labels can not only retain the correctness of the target label but also reflect richer knowledge information of the non-target label. Several studies attempt to explain why knowledge distillation is effective through various views, e.g., Re-weighting [12], Privileged Information [30], and Label-smoothing [53]. Recently, Li et al. [25] argue that correctness, smooth regularization, and class discriminability in teacher’s predicted labels can make up the knowledge of the teacher’s predicted distribution, and that appropriate category discriminability can improve the effectiveness of knowledge distillation. While achieving impressive robust performance, we are curious if adversarial robustness distillation [13, 60, 61] could address the trade-off between accuracy and robustness.
@@
Generally speaking, if we want to improve both accuracy and robustness through knowledge distillation, the provided teacher model is supposed to perform well in both aspects. That is to say, the teacher model should equally provide correct guidance and rich knowledge information for clean and adversarial examples. However, due to the existing trade-off, the teacher model itself is difficult to achieve the above goals. Therefore, a reasonable approach is to decouple adversarial knowledge distillation, i.e., using two teachers that are good at accuracy and robustness respectively (they can be called the clean teacher and the robust teacher) to guide the student in different knowledge types. Under this divide-and-conquer strategy, the student model guided by multiple teacher models can enhance its performance in both aspects theoretically.
@@
However, the actual implementation of this idea is challenging. Owing to the intrinsic trade-off between accuracy and robustness, the clean teacher and robust teacher push the student model in opposite directions. This is obviously different from the current multi-teacher knowledge distillation methods [37, 52, 29], where their optimization goal of different teachers is consistent, i.e., all teachers push the student towards the good clean accuracy. The trade-off in our task leads to a difficult optimization mainly in two aspects: (1) How to make two teacher models have similar knowledge scales. When one teacher has too much knowledge scale, the student will be forced to obtain more knowledge from this teacher and relatively less knowledge from another, which directly results in an imbalance between accuracy and robustness. However, finding a clean teacher and a robust teacher with similar knowledge scales is not easy because the different network architectures and training methods will lead to a natural gap in knowledge scales between different teachers. Thus, we need a metric to accurately measure and balance teachers’ knowledge scales. (2) How to balance the student’s learning ability from different types of teachers’ knowledge. In our multi-teacher setting, because two opposite optimization directions exist in adversarial robustness distillation, the student may over-fit an ability that is easier to learn but ignores another due to different learning difficulties. Only the cooperation of the student and teachers can help the student obtain both strong accuracy and robustness. Thus, we need an adaptive balance mechanism to adjust the learning process.
@@
Based on the above discussions, in this paper, we propose the Balanced Multi-Teacher Adversarial Robustness Distillation (B-MTARD) to enhance both accuracy and robustness at the same time. During Adversarial Training, a clean teacher is to instruct the student for handling clean examples and a strong robust teacher is to instruct the student for handling adversarial examples. To meet the two challenges above, the novel Entropy-Based Balance and Normalization Loss Balance strategies are presented to achieve a balanced state between accuracy and robustness.
@@
For the second challenge, to maintain the relative equality of the student’s learning speed from multiple teachers, we propose a Normalization Loss Balance algorithm to control loss weight and balance the influence between the clean teacher and the robust teacher. Specifically, inspired by [7], we design a relative loss to measure the knowledge proportion that the student learns from different teachers. The different loss weights for the update of student parameters are dynamically adjusted to keep the relative loss consistent during the training process. Compared with the initial state, the student model will be forced to learn relatively equal knowledge scales from all the teacher models. Our code is available at https://github.com/zhaoshiji123/MTARD-extension.
@@
  • We propose a novel framework: Balanced Multi-Teacher Adversarial Robustness Distillation (B-MTARD). B-MTARD decouples the adversarial knowledge distillation to obtain a balance between accuracy and robustness. We apply a clean teacher and a robust teacher to adaptively bring both clean and robust knowledge to the student.
@@
  • We empirically verify the effectiveness of B-MTARD in improving performance. The Weighted Robust Accuracy (a measure to evaluate both clean and robust accuracy) of our B-MTARD trained models improves significantly against a variety of attacks compared to state-of-the-art Adversarial Training and knowledge distillation methods. Besides, we show B-MTARD can achieve an obvious improvement compared to MTARD proposed in our conference version.
This journal paper is an extended version of our ECCV paper (MTARD) [59]. Compared with the conference version, we have made significant improvements and extensions in this version in the following aspects: (1) At the idea level, we discuss the trade-off problem in a more comprehensive view from both teachers and the student, while the conference version only considers student’s view. We point out that the difference in two teachers’ knowledge scales will affect the trade-off between accuracy and robustness, and should formulate this factor into our method (the first challenge in Section 1). (2) At the method level, a multiple teacher’s knowledge adjustment mechanism named the Entropy-Based Balance algorithm is designed, where we first give the definition and metric for the knowledge scale and then present the balancing method (Section 3.2). (3) At the experimental level, we give a comprehensive comparison between B-MTARD with our conference version MTARD to show the advantage, and also additionally conduct experiments on the Tiny-ImageNet dataset and compare our B-MTARD with more SOTA methods against more advanced attacks. In addition, we give more ablation studies to comprehensively test our method (Section 4).
The rest of the paper is organized as follows: Related work is given in Section 2. Section 3 introduces the details of our B-MTARD. The experiments are conducted in Section 4, and the conclusion is given in Section 5.
@@ 2.1 Adversarial Attacks
Since Szegedy et al. [39] propose that adversarial examples can mislead the deep neural network, lots of effective adversarial attack methods, such as the Fast Gradient Sign Method (FGSM) [14], Projected Gradient Descent Attack (PGD) [31], and Carlini and Wagner Attack (CW) [4] are proposed. Attack methods can be divided into white-box attacks and black-box attacks. White-box attacks know all the parameter information of the target model when generating adversarial examples, and black-box attacks know little or no information. In general, black-box attacks simulate the model gradient by repeatedly querying the target model (query-based attack) [2, 47, 44, 45, 46, 43, 27] or searching for an alternative model similar to the target model (transfer-based attack) [11, 21]. Recently, a strong adversarial attack method named AutoAttack (AA) [10] is proposed, which consists of four attack methods, including Auto-PGD (APGD), Difference of Logits Ratio (DLR) attack, FAB-Attack [9], and the black-box Square Attack [2].
@@ 2.2 Adversarial Training
To alleviate the trade-off between accuracy and robustness, some methods are proposed from different perspectives, e.g., optimization [55, 42, 32, 50] and extra data [1, 5, 19]. From the perspective of optimization, Zhang et al. [55] attempt to reduce the gap between accuracy and robustness by minimizing a Kullback–Leibler (KL) divergence loss (TRADES). Wang et al. [42] further improve performance through Misclassification-Aware Adversarial Training. Stutz et al. [38] claim that manifold analysis can be helpful in achieving accuracy and robustness. Yang et al. [50] argue that the trade-off can be mitigated by optimizing the local Lipschitz functions. Pang et al. [32] propose to employ local equivariance to describe the ideal behavior of a robust model, which facilitates the reconciliation between accuracy and robustness (SCORE). From the perspective of data, [1] and [5] find that additional unlabeled data can be useful to improve both the accuracy and robustness. [19] shows that adversarial pre-training with extra data can remarkably improve accuracy and adversarial robustness.
Different from the previous methods, we try to alleviate the trade-off by utilizing multi-teacher adversarial distillation, which applies a clean teacher and a robust teacher to bring both clean and robust knowledge to the student.
@@ 2.3 Knowledge Distillation
where $S$ represents the student model with weight $\theta_{S}$, $T$ represents the teacher model. $KL$ is Kullback–Leibler divergence loss, $CE$ represents the cross-entropy loss. $\tau$ is a temperature constant combined with softmax operation, $\alpha$ is a weight hyper-parameter.
The original knowledge distillation [20] attempts to improve the student’s performance on clean examples, and pre-define a high temperature until the teacher produces suitably soft labels, while the same high temperature is applied to train the student to match these soft labels, which allows the student to better acquire knowledge from a single type of teacher. Our method (with our Entropy-Based Balance Algorithm) is to balance the student’s performance between accuracy and robustness, and we adaptively adjust the temperature for different types of teachers to balance their knowledge scales, which allows the student to equally acquire knowledge from different types of teachers.
@@
Some studies also exist on multi-teacher knowledge distillation, and are designed from different views: including response-based knowledge [23, 52], feature-based knowledge [29, 57], and relation-based knowledge [51, 48]. Previous methods often use multiple teachers for knowledge distillation for the same examples, and those teachers tend to optimize the student in similar directions. Different from previous methods, our multi-teacher adversarial robustness distillation targets two different kinds of examples. Since adversarial examples are generated to mislead the model, there is a huge gap between the optimization direction of clean examples and adversarial examples, which exists great difficulties compared with previous research.
Figure 1: The framework of our Balanced Multi-Teacher Adversarial Robustness Distillation (B-MTARD). In the training process, we first generate adversarial examples from the student. Then, with the guidance of the clean teacher and the robust teacher, the student is trained on clean examples and adversarial examples, respectively. Here, we apply the Entropy-Based Balance algorithm to adjust the teacher’s knowledge scales; In addition, we use the Normalization Loss Balance algorithm to balance the student’s knowledge learning speed from different teachers.
@@ 2.4 Adversarial Robustness Distillation
To defend against the adversarial attack, defense distillation [33] is proposed to apply the traditional knowledge distillation. However, this method is considered as “Obfuscated Gradients” [3] and can be broken by CW attack [4]. Then a series of methods [61, 13, 60] apply knowledge distillation to further enhance the robustness based on Adversarial Training [31]. ARD [13] first proposes that using a strong robust model as the teacher model in the framework of Adversarial Training can achieve pretty robustness. IAD [60] performs adversarial knowledge distillation by compositing with unreliable teacher guidance and student introspection. RSLAD [61] uses the soft labels generated by the robust teacher model to produce adversarial examples, and further utilizes the robust labels to guide the training process of both clean and adversarial examples, which can effectively improve the robustness of the student. Fair-ARD [49] is proposed to explore the fairness issues in ARD and enhance robust fairness by re-weighting different classes. Inspired by this paper, ABSLD [58] tries to obtain a model with robust fairness by reducing the student’s class-wise error risk gap and adjusting the class-wise smoothness degree of teacher labels via re-temperating different classes.
The above methods focus on improving the robustness or fairness of ARD. In our method, we consider both accuracy and robustness through clean and robust teachers to further mitigate the trade-off, which exists an obvious difference compared with other methods.
@@ 3 Methodology
In this section, we propose B-MTARD to enhance both accuracy and robustness by applying the clean teacher and the robust teacher to guide the student. From the teachers’ teaching view, the Entropy-Based Balance algorithm is proposed to balance the teacher’s knowledge scale. From the student’s learning view, the Normalization Loss Balance algorithm is designed to balance the student’s learning speeds from different teachers.
3.1 Balanced Multi-Teacher Adversarial Distillation
The training process in B-MTARD can be considered as a min-max optimization based on Adversarial Training. The whole framework can be viewed in Fig. 1. In the inner maximization, the adversarial examples are generated by the student based on the clean examples. In the outer minimization, the clean teacher and the robust teacher generate predicted labels to guide the student training process toward the clean examples and the adversarial examples, respectively. The optimization of the basic B-MTARD is defined as follows: